OpenStudy (anonymous):

Marks: 5 [10.01] Which is the simplified form of 12 times f to the sixth power times g all over 16 times g squared.? Choose one answer. a. f to the sixth power over 4 times g squared. b. 3 times f to the sixth power all over 4 times g squared. c. 3 times f to the sixth power all over 4 times g. d. f to the sixth power over 4 times g.

OpenStudy (richyw):

\[\frac{12f^6g}{16g^2}\]there is no addition anywhere here so this is easy. Just divide! \[\frac{12}{16}=\frac{3}{4}\]\[\frac{g}{g^2}=g^{-1}\]so you get\[\frac{3f^6}{4g}\]